OpenStudy (anonymous):

Find the exact value of cos(-7pi/4)

OpenStudy (anonymous):

pi meaning 3.14...?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so that would be -7 times 3.14/4

OpenStudy (anonymous):

okay... so first times -7 by 3.14

OpenStudy (anonymous):

which is -21.98

OpenStudy (anonymous):

next divid -21.98 by 4

OpenStudy (anonymous):

which gives you -5.495

OpenStudy (jdoe0001):

\(\cfrac{7}{4}\implies \cfrac{4+3}{4}\implies \cfrac{\cancel{4}}{\cancel{4}}+\cfrac{3}{4}\implies 1+\cfrac{3}{4} \\ \quad \\ \cfrac{7\pi }{4}\implies \cfrac{4\pi +3\pi }{4}\implies \cfrac{\cancel{4}\pi }{\cancel{4}}+\cfrac{3\pi }{4}\implies \pi +\cfrac{3\pi }{4} \\ \quad \\ -\cfrac{7\pi }{4}\implies -\cfrac{4\pi +3\pi }{4}\implies -\cfrac{\cancel{4}\pi }{\cancel{4}}-\cfrac{3\pi }{4}\implies -\pi -\cfrac{3\pi }{4}\)

OpenStudy (anonymous):

hope that helps :)

OpenStudy (jdoe0001):

bear in mind that, when using a fraction of \(\pi\), is still just a fraction, so, you can siimply take the "wholes" from it, like on any fraction

OpenStudy (anonymous):

@Wow2007 was that correct?

OpenStudy (anonymous):

No, because they have to be in like fraction form

OpenStudy (anonymous):

No, because they have to be in like fraction form

OpenStudy (jdoe0001):

so... what quadrant did you end up at anyway?

OpenStudy (anonymous):

Answer A is (sqrt)2/2 Answer B is 1/2 Answer C is (sqrt)3/2 Answer D is -1/2

OpenStudy (anonymous):

I don't know, I have literally no math skills at all, I've always been way behind in math

OpenStudy (jdoe0001):

well... have you covered negative angles yet?

OpenStudy (anonymous):

Not really, these lessons are kind of everywhere. like the test before this was about surveys :/

OpenStudy (anonymous):

The book helps absolutely nothing, the only thing that is in it are definitions and examples

OpenStudy (jdoe0001):

so... if you haven't covered negative angles... how are you expected to answer it? I mean, I could ask you for my car keys only if I already know you know how to find them otherwise, it would just make no sense I mean, I could ask you just now do you know where my car keys are?

OpenStudy (anonymous):

*sigh* I know, and like you have free tutoring but its literally near impossible to get ahold of a tutor and then they want you to spend like a month learning about all kinds of things and only like 5% of it is relevant.

OpenStudy (anonymous):

I only have like 5 classes left (this is my last math) and i keep getting stuck on my math classes and it takes forever

OpenStudy (jdoe0001):

well, unless you cover the material first, it won't make much sense

OpenStudy (jdoe0001):

http://www.mathopenref.com/trigany.html <--- drag the corner over, notice when it turns negative